### v1

# Solid lines in axial algebras of Jordan type half and Jordan algebras

Abstract.
Primitive axial algebras of Jordan type $\eta$ were introduced in 2015 by Hall, Rehren and Shpectorov. These algebras are not characterized by polynomial identities. Rather, they are required to be generated by primitive idempotents whose multiplication is diagonalizable with eigenvalues $0,1,\eta$, and its eigenvectors multiply following specific fusion rules.
Based on the concept of solid subalgebras, introduced very recently by Gorshkov, Shpectorov and Staroletov, we provide a criterion for when a primitive axial algebra of Jordan type with parameter $\eta=\frac{1}{2}$ is a Jordan algebra. We also prove that this is equivalent to associators being derivations of the algebra.
Key words and phrases:
axial algebras, Jordan algebras, solid subalgebras, non-associative algebras
2020 Mathematics Subject Classification:
20B25, 20B27, 17A36, 17C27, 17D99
1. Introduction
Axial algebras were first introduced in 2015 ([6]), motivated by the classification of $2$-generated Norton-Sakuma algebras appearing in the Griess algebras of OZ vertex operator algebras ([17]). In [6], the axial algebras of Jordan type $\eta\neq\tfrac{1}{2}$ were shown to be quotients of Matsuo algebras (Definition 2.6), but the case $\eta=\tfrac{1}{2}$ has still not been classified.
In a sense, axial algebras provide an alternative to defining classes of non-associative algebras by polynomial identities, as is usually done. This provides a different framework to identify certain algebras. Especially from the viewpoint of finite group theory and sporadic groups, these algebras are more interesting. When a class of algebras is defined by polynomial laws, their automorphism groups tend to be positive-dimensional, with a very large identity component. For axial algebras of Jordan type $\eta\neq\tfrac{1}{2}$, the automorphism group is always finite, and more specifically a $3$-transposition group. This gives a way of studying $3$-transposition groups by studying axial algebras of Jordan type, see for example [2, Corollary 2.8].
In [14, Conjecture 4.3] McInroy and Shpectorov conjecture that the connected (in the sense of [14, Section 3.2]) Jordan type $\eta=\tfrac{1}{2}$ algebras are either Jordan algebras, or quotients of Matsuo algebras, certain algebras that are constructed from $3$-transposition groups. For $2$-generated and $3$-generated axial algebras of Jordan type, this has been shown to be true ([5]), and these are in fact all Jordan algebras. Very recently, Tom De Medts, Louis Rowen and Yoav Segev classified the $4$-generated case in [2], though it is still unclear whether these $4$-generated axial algebras are either Jordan algebras or quotients of Matsuo algebras.
In an effort to tackle this problem, Sergey Shpectorov introduced the idea of solid subalgebras, subalgebras of axial algebras generated by two axes for which all primitive idempotents of this subalgebra are also Jordan type $\eta=\tfrac{1}{2}$ axes of the entire algebra. The first written appearance we could find of these is [14], but they were defined in [4]. Sometimes, $2$-generated subalgebras are called lines, since $2$-generated subalgebras in a Matsuo algebra correspond to lines in the Fischer space of the underlying $3$-transposition group. In Jordan algebras, any $2$-generated subalgebra is solid (Proposition 2.8), and this is not always the case for Matsuo algebras. However, examples of non-Jordan Matsuo algebras containing both solid lines and non-solid lines do exist, and an example was found by Ilya Gorshkov and Alexei Staroletov (see [4, Example 7.2]), namely $M(3^{3}\colon S_{4})$, the Matsuo algebra constructed from the $3$ transposition group $3^{3}\colon S_{4}$ as in Definition 2.6.
In this paper, we will prove two main theorems. First, we reprove [4, Theorem 1] and extend [4, Theorem 2] to positive characteristic. This new proof does not rely on the classification of $3$-generated axial algebras of Jordan type $\eta=\tfrac{1}{2}$. The result is summarized in Theorem 1.1.
Theorem 1.1.
Let $(A,X)$ be a primitive axial algebra of Jordan type $\eta=\tfrac{1}{2}$ over a field $\mathbb{F}$ with $\operatorname{char}\mathbb{F}\neq 2$. Denote by (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) the unique Frobenius form on $A$. Given $a,b\in X$, the subalgebra $\langle\langle a,b\rangle\rangle$ is solid whenever $(a,b)\neq\frac{1}{4}$ or $\langle\langle a,b\rangle\rangle$ is not $3$-dimensional.
Second, we prove Theorem 1.2, which was already conjectured in [4, Conjecture 2], using inner derivations or associators $D_{a,b}\colon A\to A\colon x\mapsto a(xb)-b(xa)$ for $a,b\in A$, a well known concept in Jordan theory (see e.g. [13, p. 35]). To do this, we need to delve into the theory of almost Jordan rings, a less well known concept that was first studied by Marshall Osborn [16, 10]. It turns out that an $\mathbb{F}$-algebra $A$ (with $\operatorname{char}\mathbb{F}\neq 2$) is almost Jordan precisely when all associators $D_{a,b}$ are derivations of the algebra $A$. Now, it is a peculiar phenomenon that existence of sufficiently many idempotents in an almost Jordan algebra will force this algebra to be a Jordan algebra. See [16, Theorem 2] and Proposition 6.4 for examples. This allows us to prove the following statement, which is the main theorem of this article.
Theorem 1.2.
Let $(A,X)$ be a primitive axial algebra of Jordan type $\eta=\tfrac{1}{2}$ over a field $\mathbb{F}$ with $\operatorname{char}\mathbb{F}\neq 2$. If the subalgebra $\langle\langle a,b\rangle\rangle$ is solid for all $a,b\in X$, then $A$ is a Jordan algebra.
Outline of the paper
In Section 2, we define axial algebras, and give the background necessary to understand the main results in this paper. Next, in Section 3 we recall the structure theory of $2$-generated primitive axial algebras of Jordan type $\tfrac{1}{2}$, and the structure of their variety of primitive idempotents from [4, 6].
Afterwards, in Section 4, we prove Theorem 1.1. Given a $2$-generated subalgebra $\langle\langle a,b\rangle\rangle$ of a primitive axial algebra of Jordan type half, we define a class of polynomials $P_{x,y},Q_{x}\in A[\lambda]$ of small degree such that axes in $\langle\langle a,b\rangle\rangle$ and roots of the polynomials $P_{x,y},Q_{x}$ correspond to each other. Using this, we can prove that once we have enough axes in $\langle\langle a,b\rangle\rangle$ (where “enough” depends on the degree of $P_{x,y},Q_{x}$ and thus the isomorphism type of $\langle\langle a,b\rangle\rangle$), in fact all primitive idempotents of $\langle\langle a,b\rangle\rangle$ are axes.
Using the same polynomials $P_{x,y},Q_{x}$ and some base change arguments, we are able to characterize solid lines using the maps $D_{a,b}$ (Definition 2.12) for $a,b\in X$ in Section 5. It turns out that when the map $D_{a,b}$ is a derivation of the algebra $A$, all roots of the polynomials $P_{x,y},Q_{x}$ also have to be roots of their formal derivatives, and thus all roots have at least multiplicity $2$. In the same way as in Section 4, this will imply $\langle\langle a,b\rangle\rangle$ is solid.
Lastly, in Section 6, we use this new characterization of solid lines to prove Theorem 1.2. In fact, algebras $A$ where the maps $D_{x,y}$ are derivations for all $D_{x,y}$ have been studied in the 1960s by Marshall Osborn [16], and were later called almost Jordan algebras [10]. They are closely related to Jordan algebras, but do not necessarily satisfy power-associativity. However, in [16], Osborn proved that the existence of many idempotents in an almost Jordan algebra force it to satisfy power-associativity, and thus be an actual Jordan algebra. In a similar vein, we prove Proposition 6.4, which states that an almost Jordan algebra linearly spanned by idempotents is actually a Jordan algebra. We then use some induction arguments to strengthen this result a bit further, and finally prove our main theorem, Theorem 1.2.
Acknowledgements
The author is supported by the FWO PhD mandate 1172422N. The author is grateful to his supervisor Tom De Medts for his guidance, and to Sergey Shpectorov for some very insightful discussions.
2. Preliminaries
The first ingredient we need to define axial algebras, is the concept of a fusion law.
Definition 2.1.
A fusion law $(\mathcal{F},\star)$ is a set $\mathcal{F}$ with a map :×$\star\colon\mathcal{F}\times\mathcal{F}\to 2^{\mathcal{F}}$. Here $2^{\mathcal{F}}$ denotes the set of all subsets of $\mathcal{F}$.
[…]
Probably the two most important examples of fusion laws are the Jordan fusion law $\mathcal{J}(\eta)$ and the Monster fusion law $\mathcal{M}(\alpha,\beta)$ (see Table 1). The Jordan type fusion law with $\eta=\tfrac{1}{2}$ first appeared in the structure theory of Jordan algebras (see e.g. [13, Chapter III]), while the Monster fusion law has connections to the Griess algebra, the Monster group and vertex operator algebras (see e.g. [12]). For a recent survey on these two fusion laws, see [14].
$\begin{array}[]{c || c | c | c}\star&1&0&\eta\\ \hline\cr\hline\cr 1&1&&\eta\\ \hline\cr 0&&0&\eta\\ \hline\cr\eta&\eta&\eta&0,1\par\par\end{array}\hskip 34.44434pt\begin{array}[]{c || c | c | c | c}\star&1&0&\alpha&\beta\\ \hline\cr\hline\cr 1&1&&\alpha&\beta\\ \hline\cr 0&&0&\alpha&\beta\\ \hline\cr\alpha&\alpha&\alpha&1,0&\beta\\ \hline\cr\beta&\beta&\beta&\beta&1,0,\alpha\par\par\end{array}$
Table 1. The Jordan fusion law $\mathcal{J}(\eta)$ and the Monster fusion law $\mathcal{M}(\alpha,\beta)$
Using these fusion laws, we can now define axial algebras.
Definition 2.2.
Let $\mathcal{F}$ be a fusion law and $A$ a commutative non-associative $R$-algebra, where $R$ is a unital, associative, commutative ring.
  1. (i)
    For $a\in A$, let $\operatorname{ad}(a)$ denote the endomorphism of $A$ defined by multiplying elements with $a$. If $\lambda\in R$ is an eigenvalue of $\operatorname{ad}(a)$, the $\lambda$-eigenspace will be denoted by $A_{\lambda}(a)$.
  2. (ii)
    An idempotent $a\in A$ is an $\mathcal{F}$-axis if $\operatorname{ad}(a)$ is semisimple, $\operatorname{Spec}(\operatorname{ad}(a))\subseteq\mathcal{F}$ and for all $\lambda,\mu\in\operatorname{Spec}(\operatorname{ad}(a))$:
    $A_{\lambda}(a)A_{\mu}(a)\subseteq\bigoplus_{\nu\in\lambda\star\mu}A_{\nu}(a).$
    An $\mathcal{F}$-axis is primitive if $A_{1}(a)=\langle a\rangle$.
  3. (iii)
    $(A,X)$ is a (primitive) $\mathcal{F}$-axial algebra if $X\subset A$ is a set of (primitive) $\mathcal{F}$-axes that generate $A$. If $|X|=k\in\mathbb{N}$, we will call $(A,X)$ a $k$-generated $\mathcal{F}$-axial algebra.
For a fusion law we will always assume $1\in\mathcal{F}$, as axes are idempotents. We will often also write $A$ instead of $(A,X)$ for convenience. In what follows, we are interested in primitive $\mathcal{J}(\tfrac{1}{2})$-axial algebras. We will call these algebras primitive axial algebras of Jordan type (one) half to avoid using heavy notation.
Note that we defined axial algebras over rings instead of over fields, which is more common. We only do this because we want to make some arguments which require base change, while still retaining the same language. See for example Proposition 5.1.
Properties of fusion laws can lead to properties of axial algebras, as is the case for the Seress property.
Definition 2.3 (Seress property).
A fusion law $\mathcal{F}$ is called Seress if both $0,1\in\mathcal{F}$ and for every $\lambda\in\mathcal{F}$ we have $1\star\lambda,0\star\lambda\subseteq\{\lambda\}$.
Lemma 2.4 (Seress lemma).
Let $A$ be an axial algebra with fusion law $\mathcal{F}$. If $\mathcal{F}$ is Seress, then every axis $a$ associates with $A_{1}(a)\oplus A_{0}(a)$.
Proof.
See [7, Proposition 3.9]. ∎
What makes the Jordan and Monster fusion law special and interesting to study is their connection to group theory. These are examples of $\mathbb{Z}/2\mathbb{Z}$-graded fusion laws, so we can use the axes in these algebras to explicitly construct automorphisms of the algebra.
Definition 2.5 ([3, Definition 2.5]).
  1. (i)
    A fusion law $(\mathcal{F},\star)$ is called $\mathbb{Z}/2\mathbb{Z}$-graded if $\mathcal{F}$ can be partitioned into two subsets $\mathcal{F}_{+}$ and $\mathcal{F}_{-}$ such that
[…]
  2. (ii)
    Let $(A,X)$ be a $(\mathcal{F},\star)$-axial algebra for some $\mathbb{Z}/2\mathbb{Z}$-graded fusion law $(\mathcal{F},\star)$. We associate to each $(\mathcal{F},\star)$-axis $a$ of $A$ a Miyamoto involution $\tau_{a}\in\mathrm{Aut}(A)$ defined by linearly extending
    $x^{\tau_{a}}=\begin{cases}\phantom{-}x&\text{if $x\in A_{\mathcal{F}_{+}}(a)$},\\ -x&\text{if $x\in A_{\mathcal{F}_{-}}(a)$}.\end{cases}$
    Because of the $\mathbb{Z}/2\mathbb{Z}$-grading of the fusion rule, these maps define automorphisms of $A$.
  3. (iii)
    We call the subgroup $\langle\tau_{e}\mid e\in X\rangle\leq\mathrm{Aut}(A)$ the Miyamoto group of the axial algebra $(A,X)$, and we denote it by $\operatorname{Miy}(A,X)$.
The most basic example of Jordan type axial algebras are so-called Matsuo algebras. They arise from $3$-transposition groups. For more information on $3$-transposition groups, see [1, 9].
Definition 2.6.
  1. (i)
    A $3$-transposition group $(G,D)$ is a group $G$ generated by a conjugacy class of order $2$ elements $D$ such that the product of any two elements in $D$ has order at most $3$.
  2. (ii)
    Given a field $\mathbb{F}$, a $3$-transposition group $(G,D)$ and $\eta\in\mathbb{F}$, we write $M_{\eta}(\mathbb{F},(G,D))$ for the algebra $\mathbb{F}D$ with multiplication
    $a\cdot b\begin{cases}a&\text{ if }a=b,\\ 0&\text{ if }o(ab)=2,\\ \tfrac{\eta}{2}(a+b-a^{b})&\text{ if }o(ab)=3.\end{cases}$
    When all parameters are clear from context, we also write $M(G)$ for $M_{\eta}(\mathbb{F},(G,D))$.
  3. (iii)
    We call direct sums of the algebras above Matsuo algebras.
It is easy to check that Matsuo algebras $(M,D)$ with parameter $\eta$ are primitive axial algebras with respect to the fusion law $\mathcal{J}(\eta)$. In fact, whenever $\eta\neq\frac{1}{2}$, these are essentially the only examples of primitive $\mathcal{J}(\eta)$-axial algebras [6, Theorem 1.3]. It is also easy to see that the assignment
$D\to\operatorname{Miy}(M_{\eta}(G,D),D)\colon a\mapsto\tau_{a}$
induces an isomorphism of $G/Z(G)$ and $\operatorname{Miy}(M_{\eta}(G,D),D)$.
When $\eta=\frac{1}{2}$, the situation is much less clear. Shpectorov, Gorshkov and Staroletov introduced the notion of solidness of $2$-generated subalgebras of a primitive $\mathcal{J}(\tfrac{1}{2})$-axial algebra as a tool to help classify these algebras. For convenience, we will sometimes call $2$-generated subalgebras lines.
Definition 2.7.
Let $(A,X)$ be a primitive axial algebra of Jordan type half over a ring $R$ with $2\in R^{\times}$. For any $a\neq b\in X$ we will call the $2$-generated subalgebra $B=\langle\langle a,b\rangle\rangle$ solid if every primitive idempotent in $B$ is a Jordan type half axis for $A$.
Proposition 2.8.
Any $2$-generated subalgebra in a Jordan algebra is solid.
Proof.
Any idempotent in a Jordan algebra satisfies the Pierce decomposition [13, Section III.1, Lemma 1]. This implies that any idempotent satisfies the Jordan type half fusion law, and thus any $2$-generated subalgebra is automatically solid. ∎
The last tool that we will need is the existence of a Frobenius form, i.e. a bilinear form on $A$ that associates with the product.
Definition 2.9.
A bilinear form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on an algebra $A$ is a Frobenius form if
$(a,b\cdot c)=(a\cdot b,c)$
for all $a,b,c\in A$.
Lemma 2.10.
Given a primitive axial algebra $(A,X)$ of Jordan type half over a field $\mathbb{F}$, $\operatorname{char}\mathbb{F}\neq 2$, there always exists a unique Frobenius form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on $A$ such that $(a,a)=1$ for every $a\in X$. Moreover, (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) is invariant under $\mathrm{Aut}(A)$.
Proof.
This is [8, Theorem 4.1 and Lemma 4.3]. ∎
From here on out we will use (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) as notation for the Frobenius form of a given axial algebra $(A,X)$ of Jordan type half over a field $\mathbb{F}$. We immediately have some basic computations which will be used throughout.
Lemma 2.11.
Let $(A,X)$ be a primitive axial algebra of Jordan type half over a ring $R$ with $2\in R^{\times}$, with Frobenius form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) such that $(a,a)=1$ for all $a\in X$. Let $x\in A$ and $a\in X$ an axis. Then we have
  1. (i)
    $ax=\tfrac{1}{4}(x-x^{\tau_{a}})+(a,x)a,$
  2. (ii)
    $x^{\tau_{a}}=x+4(a,x)a-4ax,$
  3. (iii)
    $a(ax)=\frac{1}{2}(ax+(a,x)a).$
Proof.
As $a$ is an axis, we can write $x=x_{0}+x_{1}+x_{1/2}$ in a unique way as a sum of eigenvectors of $\operatorname{ad}(a)$, where $x_{\lambda}\in A_{\lambda}(a)$. Then $x^{\tau_{a}}=x_{0}+x_{1}-x_{1/2}$, so $x_{1/2}=\frac{1}{2}(x-x^{\tau_{a}})$. Moreover since $a$ is primitive, we have $x_{1}=\lambda a$ for some $\lambda\in R$, and $(a,x)=(a,x_{1})=\lambda$. Thus $x=x_{0}+(a,x)a+\frac{1}{2}(x-x^{\tau_{a}})$. This implies $ax=\frac{1}{4}(x-x^{\tau_{a}})+(a,x)a$. Clearly (i),(ii) and (iii) now follow. ∎
Lastly, we will need the associators alluded to in the introduction.
Definition 2.12.
Let $A$ be a commutative algebra, and $a,b\in A$. We define the map $D_{a,b}\colon A\to A$ by $D_{a,b}(x)=a(bx)-b(ax)$.
3. $2$-generated subalgebras
We will give a quick review of $2$-generated axial algebras of Jordan type half. For more information on this topic, we refer to [4, 6]. Any $2$-generated axial algebra of Jordan type half is a quotient of a $3$-dimensional $2$-generated axial algebra of Jordan type half. We can classify them by the value of the Frobenius form $(a,b)$, where $a$ and $b$ are generating axes of the $2$-generated axial algebra $B$ [6, Theorem 4.6]. They come in three different flavors: toric, baric and flat lines. We discuss each of them separately, and review some facts we will use.
Definition 3.1.
Let $\langle\langle a,b\rangle\rangle$ be a $2$-generated primitive axial algebra of Jordan type half, and denote by (,)(,)( , ) its unique Frobenius form from Lemma 2.10.
  1. (i)
    if $(a,b)\neq 0,1$ we call $\langle\langle a,b\rangle\rangle$ toric,
  2. (ii)
    If $(a,b)=0$ we call $\langle\langle a,b\rangle\rangle$ flat,
  3. (iii)
    if $(a,b)=1$ we call $\langle\langle a,b\rangle\rangle$ baric.
Moreover, we collectively call the flat and baric algebras unipotent.
One major distinction between toric lines and flat or baric lines is the structure of their automorphism groups. In the toric case, the identity component of the automorphism group will be a multiplicative group, or say, a torus. In case the $2$-generated algebra is baric or flat, the identity component of its automorphism group is unipotent. We will often collectively call the baric and flat cases the unipotent case. For what we will try to prove, it is necessary to be aware of this distinction, as it means the arguments given for these two cases will differ slightly in details.
3.1. Toric algebras
In this subsection, we describe toric algebras, the variety of primitive idempotents in toric algebras and the action of the Miyamoto involutions.
Lemma 3.2.
Let $B=\langle\langle a,b\rangle\rangle$ be a toric algebra over an algebraically closed field $\mathbb{F}$. Then $B$ is a $3$-dimensional simple unital Jordan algebra (with unit $u$), and there exists a basis $\{e,u,f\}$ of $B$ with $e^{2}=f^{2}=0$, $ef=\frac{1}{8}u$, $a=e+\frac{1}{2}u+f$ and $b=\mu e+\frac{1}{2}u+f$ for a certain $\mu\in\mathbb{F}^{\times}$.
Proof.
See [4, Lemma 3.7, Lemma 3.8]. Note that we can always rescale $e,f$ such that $a=e+\frac{1}{2}u+f$. ∎
Definition 3.3.
For $\lambda\in\mathbb{F}^{\times}$, we write $a_{\lambda}=\lambda e+\tfrac{1}{2}u+\lambda^{-1}f$.
Lemma 3.4.
Any primitive idempotent in the toric algebra $\langle\langle a,b\rangle\rangle$ is of the form $a_{\lambda}$.
Proof.
See [4, Lemma 3.5]. ∎
Lastly, we need to know a bit about the action of the Miyamoto involutions on the primitive idempotents. Note that every primitive idempotent satisfies the Jordan fusion law, since $\langle\langle a,b\rangle\rangle$ is Jordan.
Lemma 3.5.
Let $\tau_{a_{\mu}}$ be the Miyamoto involution corresponding to the axis $a_{\mu}$. Then
$a_{\lambda}^{\tau_{a_{\mu}}}=a_{\lambda^{-1}\mu^{2}}\text{ for all }\lambda,\mu\in\mathbb{F}^{\times}.$
Proof.
This follows from [4, Lemma 3.9]. ∎
Corollary 3.6.
Let $\langle\langle a,b\rangle\rangle$ be toric. Write $G=\langle\tau_{a},\tau_{b}\rangle$. The union of orbits $O=a^{G}\cup b^{G}$ has size $n<\infty$ if and only if $\mu$ is an $n$-th root of unity, where $\mu$ is as in Lemma 3.2.
Proof.
Note that $a=a_{1}$ and $b=a_{\mu}$. Given $k\in\mathbb{Z}$, we get that $a^{(\tau_{a}\tau_{b})^{k}}=a_{\mu^{2k}}\in O$ and $b^{(\tau_{a}\tau_{b})^{k}}=a_{\mu^{2k+1}}\in O$. It is then easy to see that $O=\{a_{\mu^{k}}\in\langle\langle a,b\rangle\rangle\mid k\in\mathbb{Z}\}$. For this set to be finite, it is both sufficient and necessary that $\mu^{n}=1$ for a certain $n\in\mathbb{N}$, and if $n$ is minimal with respect to this property, then $|O|=n$. ∎
Corollary 3.7.
Let $\langle\langle a,b\rangle\rangle$ be toric. Write $G=\langle\tau_{a},\tau_{b}\rangle$. If $O=a^{G}\cup b^{G}$ has size $3$, then $(a,b)=\tfrac{1}{4}$.
Proof.
We know from Corollary 3.6 that in this case, $a=e+\frac{1}{2}u+f,b=\mu e+\frac{1}{2}u+\mu^{2}f$, where $\mu$ is a primitive third root of unity. Using the description of the Frobenius form in [4, p.12-13], we get that $(a,b)=\frac{1}{4}(mu+mu^{2})+\frac{1}{2}$. Since $\mu$ is a primitive third root of unity, we have $\mu^{2}+\mu=-1$, and thus $(a,b)=\tfrac{1}{4}$. ∎
3.2. Flat algebras
In this subsection, we describe flat algebras, the variety of primitive idempotents in flat algebras and the action of the Miyamoto involutions.
Lemma 3.8.
Let $B=\langle\langle a,b\rangle\rangle$ be a flat algebra over a field $\mathbb{F}$. Then $B$ is Jordan and either
  1. (i)
    a $3$-dimensional algebra, denoted $\mathfrak{J}(0)$, with basis $\{a,b,v=ab\}$,
  2. (ii)
    or a $2$-dimensional quotient of this algebra, isomorphic to $\mathbb{F}\oplus\mathbb{F}$.
Proof.
See [4, Theorem 3.2(a)]. ∎
Definition 3.9.
For $\lambda\in\mathbb{F}$, we write $a_{\lambda}=a+\lambda v$ and $b_{\lambda}=b+\lambda v$ with $v=ab$.
Lemma 3.10.
Any primitive idempotent in the flat algebra $\langle\langle a,b\rangle\rangle$ is of the form $a_{\lambda}$ or $b_{\lambda}$.
Proof.
See [4, Proposition 6.6]. The calculations were made for $\operatorname{char}\mathbb{F}=0$, but remain true for $\operatorname{char}\mathbb{F}\neq 2$. ∎
Lastly, we need to know a bit about the action of the Miyamoto involutions on the primitive idempotents. Note that every primitive idempotent satisfies the Jordan fusion law, since $\langle\langle a,b\rangle\rangle$ is Jordan.
Lemma 3.11.
Let $\tau_{a_{\mu}},\tau_{b_{\mu}}$ be the Miyamoto involution corresponding to the axis $a_{\mu},b_{\mu}$ respectively. Then
$a_{\lambda}^{\tau_{a_{\mu}}}=a_{2\mu-\lambda}\text{ and }a_{\lambda}^{\tau_{b_{\mu}}}=a_{-4-2\mu-\lambda}\text{ for all }\lambda,\mu\in\mathbb{F}.$
Proof.
Note that $v=ab$ is a $\tfrac{1}{2}$-eigenvector for $a$ and $b$ and $v^{2}=0$ by [4, Theorem 3.1]. We can then compute using Lemma 2.11 that
$a_{\lambda}^{\tau_{a_{\mu}}}=a_{\lambda}+4(a_{\lambda},a_{\mu})a_{\mu}-4a_{\lambda}a_{\mu}=a_{\lambda+4\mu-2\lambda-2\mu}$
and
$a_{\lambda}^{\tau_{b_{\mu}}}=a_{\lambda}+4(a_{\lambda},b_{\mu})b_{\mu}-4a_{\lambda}b_{\mu}=a_{\lambda-4-2\lambda-2\mu}.\qed$
Corollary 3.12.
Suppose $\langle\langle a,b\rangle\rangle$ is flat and not $2$-dimensional. Write $G=\langle\tau_{a},\tau_{b}\rangle$. The orbit $O=a^{G}$ has size $n<\infty$ if and only if $02\neq\operatorname{char}\mathbb{F}=p>$ 0, and then $|O|=p$.
Proof.
Note that $a=a_{0}$ and $b=b_{0}$. Given $k\in\mathbb{Z}$, we get that $a^{(\tau_{a}\tau_{b})^{k}}=a_{-4k}\in O$. It is then easy to see that $O=\{a_{4k}\in\langle\langle a,b\rangle\rangle\mid k\in\mathbb{Z}\}$. For this set to be finite, it is both sufficient and necessary that the characteristic of the field is positive, and since $\operatorname{char}\mathbb{F}=p\neq 2$, we then have that |$|\{a_{k}|k\in\{0,\dots,p-1\}\}|=p$. ∎
3.3. Baric algebras
In this subsection, we describe baric algebras, the variety of primitive idempotents in baric algebras and the action of the Miyamoto involutions.
Lemma 3.13.
Let $B=\langle\langle a,b\rangle\rangle$ be a baric algebra over a field $\mathbb{F}$. Then $B$ is Jordan and either
  1. (i)
    a $3$-dimensional algebra, denoted $\mathfrak{J}(1)$, with basis $\{a,v=2(ab-a),v^{2}\}$,
  2. (ii)
    a $2$-dimensional quotient of this algebra, denoted $\overline{\mathfrak{J}(1)}$, with basis $\{a,v=2(ab-a)\}$,
  3. (iii)
    or a $1$-dimensional quotient.
Proof.
See [4, Theorem 3.2(b)]. ∎
We will not consider the $1$-dimensional quotient for obvious reasons.
Definition 3.14.
For $\lambda\in\mathbb{F}$, we write $a_{\lambda}=a+\lambda v+\lambda^{2}v^{2}$ with $v=2(ab-a)$.
Lemma 3.15.
Any primitive idempotent in the flat algebra $\langle\langle a,b\rangle\rangle$ is of the form $a_{\lambda}$.
Proof.
See [4, Proposition 6.4]. The calculations were made for $\operatorname{char}\mathbb{F}=0$, but remain true for $\operatorname{char}\mathbb{F}\neq 2$. ∎
Lastly, we need to know a bit about the action of the Miyamoto involutions on the primitive idempotents. Note that every primitive idempotent satisfies the Jordan fusion law, since $\langle\langle a,b\rangle\rangle$ is Jordan.
Lemma 3.16.
Let $\tau_{a_{\mu}}$ be the Miyamoto involution corresponding to the axis $a_{\mu}$. Then
$a_{\lambda}^{\tau_{a_{\mu}}}=a_{2\mu-\lambda}\text{ for all }\lambda,\mu\in\mathbb{F}.$
Proof.
Note that $v=2(ab-a)$ is a $\tfrac{1}{2}$-eigenvector for $a$ and $v^{2}$ is a 00-eigenvector for $a$ by [4, Theorem 3.1]. We can then compute using Lemma 2.11 that
$a_{\lambda}^{\tau_{a_{\mu}}}=a_{\lambda}+4(a_{\lambda},a_{\mu})a_{\mu}-4a_{\lambda}a_{\mu}=a_{\lambda+4\mu-2\lambda-2\mu}.\qed$
Corollary 3.17.
Suppose $\langle\langle a,b\rangle\rangle$ is baric and not $1$-dimensional. Write $G=\langle\tau_{a},\tau_{b}\rangle$. The orbit $O=a^{G}$ has size $n<\infty$ if and only if $02\neq\operatorname{char}\mathbb{F}=p>$ 0, and then $|O|=p$.
Proof.
Note that $a=a_{0}$ and $b=a_{1}$. Given $k\in\mathbb{Z}$, we get that $a^{(\tau_{a}\tau_{b})^{k}}=a_{2k}\in O$. It is then easy to see that $O=\{a_{2k}\in\langle\langle a,b\rangle\rangle\mid k\in\mathbb{Z}\}$. For this set to be finite, it is both sufficient and necessary that the characteristic of the field is positive, and since $\operatorname{char}\mathbb{F}=p\neq 2$, we then have that |$|\{a_{k}|k\in\{0,\dots,p-1\}\}|=p$. ∎
4. Proof of Theorem 1.1
Even though a large portion of this theorem was already proved in [4], we give an alternative proof here. This proof uses the same techniques for all types of lines, and it does not use the classification of $3$-generated primitive axial algebras of Jordan type [5]. To prove this fact, we will try to find polynomials of small degree which tell us when a subalgebra is solid (Lemma 4.1), and then prove they have many zeroes in most cases, which forces the polynomials to be identically zero. For the remainder of this section, $(A,X)$ is a primitive axial algebra of Jordan type half over a field $\mathbb{F}$ (assumed algebraically closed, without loss of generality) with $\operatorname{char}\mathbb{F}\neq 2$, and $a,b\in X$ are axes.
Lemma 4.1.
Let $(A,X)$ be a primitive axial algebra of Jordan type half over a field $\mathbb{F}$, and $a,b\in X$. the subalgebra $\langle\langle a,b\rangle\rangle$ is solid if and only if
$Q_{x}(c)=c(cx)-\frac{1}{2}(cx+(c,x)c)$
and
$P_{x,y}(c)=4(cx)(cy)-(c,y)cx-(cy)x-(c,x)cy-(cx)y-(c,xy)c+c(xy)$
are zero for all primitive idempotents $c\in\langle\langle a,b\rangle\rangle$ and $x,y\in A$. Moreover, is $c\in\langle\langle a,b\rangle\rangle$ satisfies the Jordan fusion law, then $Q_{x}(c)=P_{x,y}(c)=0$ for all $x,y\in A$.
Proof.
Given a primitive idempotent $c\in\langle\langle a,b\rangle\rangle$, let $\phi_{c}\colon A\to A$ be the map defined by $x^{\phi_{c}}=x+4(c,x)c-4cx$.
Suppose $P_{x,y}(c)=0=Q_{x}(c)$ for all $x,y\in A,c\in\langle\langle a,b\rangle\rangle$. We will show $\phi_{c}$ is an automorphism of $A$ for every primitive idempotent $c\in\langle\langle a,b\rangle\rangle$. For every $x,y\in A$ we have
$x^{\phi_{c}}y^{\phi_{c}}-(xy)^{\phi_{c}}$ $=(x+4(c,x)c-4cx)(y+4(c,y)c-4cy)-(xy+4(c,xy)c-4c(xy))$
$=4(c,y)cx-4(cy)x+4(c,x)cy+16(c,x)(c,y)c-16(c,x)c(cy)$
$\hskip 12.91663pt-4(cx)y-16(c,y)c(cx)+16(cx)(cy)-(xy+4(c,xy)c-4c(xy))$
By using that $Q_{x}(c)=Q_{y}(c)=0$ this last expression is equal to $4P_{x,y}(c)$, which is equal to zero. This means $\phi_{c}$ is an automorphism for every primitive idempotent $c\in\langle\langle a,b\rangle\rangle$. But $\{a,b\}^{\langle\phi_{c}\mid c\in\mathbb{F}\rangle}$ is equal to the variety of all primitive idempotents in $\langle\langle a,b\rangle\rangle$, which means the line is solid.
Conversely, if $c$ is an axis, then $Q_{x}(c)=0$ for every $x\in A$ by Lemma 2.11, and the calculation above then shows $P_{x,y}(c)=0$ for every $x,y\in A$, since $\phi_{c}=\tau_{c}$ is an automorphism of $A$. ∎
4.1. The toric case
We will write $a_{\lambda}=\lambda e+\tfrac{1}{2}u+\lambda^{-1}f$, where $\lambda\in\mathbb{F}^{\times}$ and $e,u,f$ as in Lemma 3.2, chosen with respect to the generating axes $a,b$ of $\langle\langle a,b\rangle\rangle$.
Definition 4.2.
Let $(A,X)$ be a primitive axial algebra of Jordan type half over $\mathbb{F}$, and $a,b\in X$ such that $\langle\langle a,b\rangle\rangle$ is toric. Let $R$ be a unital commutative associative $\mathbb{F}$-algebra, and $\lambda\in R^{\times}$ an invertible element.
  1. (i)
    We define $a_{\lambda}\coloneqq\lambda e+\tfrac{1}{2}u+\lambda^{-1}f\in\langle\langle a,b\rangle\rangle_{R}$.
  2. (ii)
    Given $x,y\in A_{R}$, we define $P^{t}_{x,y}$ and $Q^{t}_{x}$ as the maps
    $P^{t}_{x,y}\colon R^{\times}\to A\colon\lambda\mapsto P_{x,y}(a_{\lambda})$
    and
    $Q^{t}_{x}\colon R^{\times}\to A\colon\lambda\mapsto P_{x,y}(a_{\lambda}).$
Note that the maps $P^{t}_{x,y},Q^{t}_{x,y}$ are rational maps, but $\lambda^{2}P^{t}_{x,y},\lambda^{2}Q^{t}_{x,y}\in A[\lambda]$ are polynomial maps, since $a_{\lambda}=\lambda^{-1}(\lambda^{2}e+\lambda\tfrac{1}{2}u+f)$.
Proposition 4.3.
Let $(A,X)$ be a primitive axial algebra of Jordan type half and $a,b\in X$ with $(a,b)\neq 0,1,\frac{1}{4}$. Then $\langle\langle a,b\rangle\rangle$ is solid.
Proof.
Under the conditions in the statement, we know $\langle\langle a,b\rangle\rangle$ is toric, and $\langle\langle a,b\rangle\rangle$ contains at least $4$ axes by Corollary 3.7. The polynomial maps $\lambda^{2}P^{t}_{x,y},\lambda^{2}Q^{t}_{x}$ have degree less than or equal to $4$ for all $x,y\in A$. If $\langle\langle a,b\rangle\rangle$ has at least $5$ different axes, then the polynomials $\lambda^{2}P^{t}_{x,y},\lambda^{2}Q^{t}_{x}$ have at least $5$ different zeroes, and hence are identically zero. Thus, for every $\lambda\in\mathbb{F}^{\times}$ and $x,y\in A$, we have that $P^{t}_{x,y}(\lambda)=Q^{t}_{x}(\lambda)=0$, so by Lemma 4.1, the subalgebra $\langle\langle a,b\rangle\rangle$ is solid.
Now suppose $\langle\langle a,b\rangle\rangle$ has only $4$ axes. Then these have to be of the form $a_{\mu}$, with $\mu^{4}=1$ by Corollary 3.6. We write
$\lambda^{2}P^{t}_{x,y}(\lambda)=a_{0}(x,y)+a_{1}(x,y)\lambda+a_{2}(x,y)\lambda^{2}+a_{3}(x,y)\lambda^{3}+a_{4}(x,y)\lambda^{4}$
and
$\lambda^{2}Q^{t}_{x}(\lambda)=b_{0}(x)+b_{1}(x)\lambda+b_{2}(x)\lambda^{2}+b_{3}(x)\lambda^{3}+b_{4}(x)\lambda^{4}$
for every $x,y\in A$. Then one can compute
$a_{4}(x,y)$ $=4(ex)(ey)-(e,x)ey-(e,y)ex-(e,xy)e,$
$\begin{split}a_{3}(x,y)&=2(ex)(uy)+2(ux)(ey)-\tfrac{1}{2}(e,y)ux-\tfrac{1}{2}(u,y)ex\\ &\hskip 12.91663pt-(ey)x-\tfrac{1}{2}(e,x)uy-\tfrac{1}{2}(u,x)ey-(ex)y-\tfrac{1}{2}(e,xy)u-\tfrac{1}{2}(u,xy)e+e(xy),\end{split}$
$a_{0}(x,y)$ $=4(fx)(fy)-(f,x)fy-(f,y)fx-(f,xy)f,$
$b_{4}(x)$ $=e(ex)-\tfrac{1}{2}(e,x)e,$
$b_{3}(x)$ $=\tfrac{1}{2}e(ux)+\tfrac{1}{2}u(ex)-\tfrac{1}{2}ex-\tfrac{1}{4}(e,x)u-\tfrac{1}{4}(u,x)e.$
Note that $a_{3}(x,e)=2b_{4}(x)$ and $a_{3}(x,u)=4b_{3}(x)+e(ux)-u(ex)$. For all $x,y\in A$, the polynomial maps $\lambda^{2}P^{t}_{x,y},\lambda^{2}Q^{t}_{x}$ have ±$\pm 1,\pm i$ as zeroes, where $i$ is a primitive fourth root of unity by Corollaries 3.6 and 4.1. Any polynomial of degree at most $4$ which has the fourth roots of unity as roots is necessarily a scalar multiple of $x^{4}-1$. This implies that $a_{3}(x,y)=0$ for all $x,y\in A$, and thus $b_{4}(x)=0$ for all $x\in A$. This in turn implies that $Q^{t}_{x}(\lambda)=0$ for all $\lambda\in\mathbb{F}^{\times}$.
We still need to prove $a_{4}(x,y)=0$ for all $x,y\in A$. Since $a_{3}(x,u)=b_{3}(x)=0$ for all $x\in A$, we also have $e(ux)=u(ex)=\frac{1}{2}ex+\frac{1}{4}((e,x)u+(u,x)e)$ for all $x\in A$. Given $x\in A$, we compute
(1) $e(x^{\tau_{a}})=e(x+4(a,x)a-4ax)\\ =ex+2(a,x)e+\tfrac{1}{2}(a,x)u-4e(ex)-2e(ux)-4e(fx)\\ =ex+2(e,x)e+(u,x)e+2(f,x)e+\tfrac{1}{2}(e,x)u+\tfrac{1}{4}(u,x)u\\ +\tfrac{1}{2}(f,x)u-2(e,x)e-ex-\tfrac{1}{2}(e,x)u-\tfrac{1}{2}(u,x)e-4e(fx)\\ =-4e(fx)+\tfrac{1}{2}(u,x)e+2(f,x)e+\tfrac{1}{4}(u,x)u+\tfrac{1}{2}(f,x)u,$
proving that $e(fx)=-\tfrac{1}{4}e(x^{\tau_{a}})+\tfrac{1}{2}(f,x)e+\tfrac{1}{8}(u,x)e+\tfrac{1}{8}(f,x)u+\tfrac{1}{16}(u,x)u$.
Next, we show that $a_{4}(fx,y)=0$ for all $x,y\in A$:
$a_{4}(fx,y)=-a_{0}(fx,y)$ $=-4(f(fx))(fy)+(f,fx)fy+(f,y)f(fx)+(f,(fx)y)f$
$=-2(f,x)f(fy)+\tfrac{1}{2}(f,y)(f,x)f+\tfrac{1}{2}(f,y)(f,x)f$
$=-(f,y)(f,x)f+(f,y)(f,x)f$
$=0.$
On the other hand, using Equation 1, we get
$a_{4}(fx,y)$ $=4(e(fx))(ey)-(e,fx)ey-(e,y)e(fx)-(e,(fx)y)e$
$=-(e(x^{\tau_{a}}))(ey)+\tfrac{1}{2}(u,x)e(ey)+2(f,x)e(ey)+\tfrac{1}{4}(u,x)u(ey)+\tfrac{1}{2}(f,x)u(ey)-\tfrac{1}{8}(u,x)ey$
$+\tfrac{1}{4}(e,y)e(x^{\tau_{a}})-\tfrac{1}{8}(e,y)(u,x)e-\tfrac{1}{2}(e,y)(f,x)e-\tfrac{1}{16}(e,y)(u,x)u-\tfrac{1}{8}(e,y)(f,x)u$
$+\tfrac{1}{4}(e(x^{\tau_{a}}),y)e-\tfrac{1}{8}(u,x)(e,y)e-\tfrac{1}{2}(f,x)(y,e)e-\tfrac{1}{16}(u,x)(u,y)e-\tfrac{1}{8}(f,x)(u,y)e$
$=-(e(x^{\tau_{a}}))(ey)+\tfrac{1}{4}(e,y)e(x^{\tau_{a}})+\tfrac{1}{4}(e,x^{\tau_{a}}y)e+\tfrac{1}{4}(f,x)ey$
$\hskip 12.91663pt+\big{(}\tfrac{1}{2}(u,x)+2(f,x)\big{)}\big{(}e(ey)-\tfrac{1}{2}(e,y)e\big{)}$
$\hskip 12.91663pt+\big{(}\tfrac{1}{4}(u,x)+\tfrac{1}{2}(f,x)\big{)}\big{(}u(ey)-\tfrac{1}{2}ey-\tfrac{1}{4}((e,y)u+(u,y)e)\big{)}$
$=-(e(x^{\tau_{a}}))(ey)+\tfrac{1}{4}(e,y)e(x^{\tau_{a}})+\tfrac{1}{4}(e,x^{\tau_{a}}y)e+\tfrac{1}{4}(e,x^{\tau_{a}})ey$
$=-\tfrac{1}{4}a_{4}(x^{\tau_{a}},y).$
This implies $a_{4}(x^{\tau_{a}},y)=0$ for all $x,y\in A$, and since $\tau_{a}$ is a bijection, we get $a_{4}(x,y)=0$ for all $x,y\in A$. This in turn implies that $\lambda^{2}P^{t}_{x,y}$ is the zero polynomial, so $P^{t}_{x,y}(\lambda)=0$ for all $\lambda\in\mathbb{F}^{\times}$. We can now use Lemma 4.1 to show that $\langle\langle a,b\rangle\rangle$ is solid. ∎
The baric and flat cases are easier to deal with. In these cases, we usually have more axes in $\langle\langle a,b\rangle\rangle$.
4.2. The unipotent case
In [4, Section 6] it was proven that $\langle\langle a,b\rangle\rangle$ is solid for $(a,b)=0,1$ when $\operatorname{char}\mathbb{F}=0$. We will extend this result to positive characteristic.
Definition 4.4.
Let $(A,X)$ be a primitive axial algebra of Jordan type half over $\mathbb{F}$, and $a,b\in X$ such that $\langle\langle a,b\rangle\rangle$ is unipotent. Let $R$ be a unital commutative associative $\mathbb{F}$-algebra, and $\lambda\in R$.
  1. (i)
    We define $a_{\lambda}\coloneqq a+\lambda v+\lambda^{2}v^{2}\in B_{R}$, where $v=a\cdot b,2(a\cdot b-a)$ when $(a,b)=0,1$ respectively.
  2. (ii)
    Given $x,y\in A_{R}$, we define $P^{u}_{x,y}$ and $Q^{u}_{x}$ as the maps
    $P^{u}_{x,y}\colon R\to A\colon\lambda\mapsto P_{x,y}(a_{\lambda})$
    and
    $Q^{u}_{x}\colon R\to A\colon\lambda\mapsto P_{x,y}(a_{\lambda}).$
Before we state the following proposition, recall the notations for the different possible $2$-generated algebras from Lemmas 3.8 and 3.13.
Proposition 4.5.
Let $(A,X)$ be a primitive axial algebra of Jordan type half over a field $\mathbb{F}$ and $a,b\in X$ with $\langle\langle a,b\rangle\rangle\not\cong\mathfrak{J}(1)$, or $\operatorname{char}\mathbb{F}\neq 3$ and $\langle\langle a,b\rangle\rangle\cong\mathfrak{J}(1)$. Then $\langle\langle a,b\rangle\rangle$ is solid.
Proof.
Note that $P^{u}_{x,y}$ and $Q^{u}_{x}$ are always polynomial maps.
Suppose first that $\langle\langle a,b\rangle\rangle\not\cong\mathfrak{J}(1)$. If $\langle\langle a,b\rangle\rangle$ is isomorphic to $\mathbb{F}\oplus\mathbb{F}$, the proposition is clearly true since the only primitive idempotents in $\langle\langle a,b\rangle\rangle$ are $a$ and $b$. So suppose $\langle\langle a,b\rangle\rangle\cong\mathfrak{J}(0)$ or $\overline{\mathfrak{J}(1)}$. Then the subalgebra $\langle\langle a,b\rangle\rangle$ contains at least $3$ different axes $a_{\lambda}$ by Corollaries 3.12 and 3.17, so the polynomial equations $P^{u}_{x,y}(\lambda)$ and $Q^{u}_{x}(\lambda)$ from Lemma 4.1 have at least $3$ different zeroes for $x,y\in A$. However, $\deg P^{u}_{x,y},\deg Q^{u}_{x}\leq 2$, so this is only possible when $P^{u}_{x,y},Q^{u}_{x}$ are identically zero for every $x,y\in A$. For flat lines, we also have that $P_{x,y}(b-\lambda v),Q_{x}(b-\lambda v)$ are zero for all $x,y\in A,\lambda\in\mathbb{F}$ by symmetry. The result now follows from Lemma 4.1.
If $\langle\langle a,b\rangle\rangle\cong\mathfrak{J}(1)$ and $\operatorname{char}\mathbb{F}\neq 3$, then the polynomial equations $P^{u}_{x,y}(\lambda)$ and $Q^{u}_{x}(\lambda)$ from Lemma 4.1 have at least $5$ different zeroes for $x,y\in A$ by Lemmas 4.1 and 3.17, and $\deg P^{u}_{x,y},\deg Q^{u}_{x}\leq 4$. Then the result again follows from Lemma 4.1. ∎
Over fields of characteristic $3$, subalgebras isomorphic to $\mathfrak{J}(1)$ are not always solid! We can find Matsuo algebras where not every line is solid, the smallest example being $M(W(D_{4}))$. It turns out this characteristic is the only problematic one. We now establish Theorem 1.1.
Theorem 4.6.
Let $(A,X)$ be a primitive axial algebra of Jordan type one half over a field $\mathbb{F}$ with $\operatorname{char}\mathbb{F}\neq 2$. Denote by (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) the unique Frobenius form on $A$. Given $a,b\in X$, the $\langle\langle a,b\rangle\rangle$ is solid whenever $(a,b)\neq\frac{1}{4}$ or $\langle\langle a,b\rangle\rangle$ is not $3$-dimensional.
Proof.
This follows from Propositions 4.3 and 4.5. ∎
5. Derivations
In fact, in the previous section we have proved more than just the fact that lines are solid. Because we proved certain polynomials are equal to the zero polynomial, we actually have the following. From here onwards, ideas from the theory of group schemes are used, but since we only use basic techniques, we do not use the technical language. For an introduction to this theory, see [11, 15].
Proposition 5.1.
Given $a,b\in X$ such that $\langle\langle a,b\rangle\rangle$ is solid and $\lambda\in R$ (resp. $R^{\times}$ when $\langle\langle a,b\rangle\rangle$ is toric), where $R$ is a (unital, commutative, associative) $\mathbb{F}$-algebra. Then $a_{\lambda}$ is an axis of $R\otimes_{\mathbb{F}}A$.
Proof.
Let $R$ be a unital, commutative associative $\mathbb{F}$-algebra. The polynomials or rational maps $P^{t}_{x,y},Q^{t}_{x},P^{u}_{x,y},Q^{u}_{x}$ are all zero maps, and remain so after base change. We can then use the same argument as in Lemma 4.1 to show $a_{\lambda}$ is an axis. ∎
In the next proposition, we let $\mathbb{F}[\varepsilon]$ be the algebra of dual numbers, i.e. $\mathbb{F}[\varepsilon]\coloneqq\{a+\varepsilon b\mid a,b\in\mathbb{F}\}$ with $\varepsilon^{2}=0$. The following proposition is again well known, and a common way to construct the Lie algebra of automorphism group schemes.
Proposition 5.2.
Let $A$ be an $\mathbb{F}$-algebra, and $\phi$ a map of $A\otimes_{\mathbb{F}}\mathbb{F}[\varepsilon]$ such that there exists $f\colon A\to A$ with $\phi(x)=x+\varepsilon f(x)$ for all $x\in A$. Then $f$ is a derivation of $A$ if and only if $\phi$ is an automorphism of $A\otimes_{\mathbb{F}}\mathbb{F}[\varepsilon]$.
Proof.
For any $x,y\in A\subset A\otimes_{\mathbb{F}}\mathbb{F}[\varepsilon]$ we have $(xy)^{\phi}=x^{\phi}y^{\phi}=xy+\varepsilon(f(x)y+xf(y))$ if and only if $f(x)y+xf(y)=f(xy)$. Since every element of $A\otimes_{\mathbb{F}}{\mathbb{F}[\varepsilon]}$ is a $\mathbb{F}[\varepsilon]$-linear combination of elements in $A$, this implies $f$ is a derivation if and only if $\phi$ is an automorphism of $A\otimes_{\mathbb{F}}\mathbb{F}[\varepsilon]$. ∎
Corollary 5.3.
If $\langle\langle a,b\rangle\rangle$ is solid, then $D_{a,b}$ is a derivation.
Proof.
Given $a,b\in X$ such that $B=\langle\langle a,b\rangle\rangle$ is solid. By Proposition 5.1, we know $\tau_{a_{\lambda}}$ is an automorphism of $A\otimes_{\mathbb{F}}\mathbb{F}[\varepsilon]$ for $\lambda\in\mathbb{F}[\varepsilon]$ or $\mathbb{F}[\varepsilon]^{\times}$, where $\mathbb{F}[\varepsilon]$ is the algebra of dual numbers. If $\langle\langle a,b\rangle\rangle$ is toric, then we consider $\lambda=1+\varepsilon$, and $a_{\lambda}=a+\varepsilon(e-f)$. If $\langle\langle a,b\rangle\rangle$ is unipotent, then we consider $\lambda=\varepsilon$. If $\langle\langle a,b\rangle\rangle$ is flat, then $a_{\lambda}=a+\varepsilon ab$ and if $\langle\langle a,b\rangle\rangle$ is baric, then $a_{\lambda}=a+\varepsilon(b-a+2\sigma)$. In all three cases, we have $a_{\lambda}=a+\varepsilon v$ where $v$ is a $\frac{1}{2}$-eigenvector of $a$. We compute
$a(vx)$ $=\frac{1}{4}(vx+vx^{\tau_{a}})+\frac{1}{2}(v,x)a$
$=\frac{1}{4}(vx+vx+2(a,x)v-4v(ax))+\frac{1}{2}(v,x)a$
$=-v(ax)+\frac{1}{2}(vx+(v,x)a+(a,x)v),$
so $[L_{a},L_{v}](x)=-2v(ax)+\frac{1}{2}(vx+(v,x)a+(a,x)v)$. We will compute the action of $\rho=\tau_{a}\tau_{a_{\lambda}}$ on $A$, where $\lambda\in\mathbb{F}[\varepsilon]$ is chosen as above. Given $x\in A\subset A\otimes_{\mathbb{F}}\mathbb{F}[\varepsilon]$, we obtain
$x^{\rho}$ $=x^{\tau_{a}}+4(a+\varepsilon v,x^{\tau_{a}})(a+\varepsilon v)-4(a+\varepsilon v)x^{\tau_{a}}$
$=x^{\tau_{a}}+4(a,x^{\tau_{a}})a-4ax^{\tau_{a}}$
$\hskip 12.91663pt+4\varepsilon(-(v,x)a+(a,x)v-v(x+4(a,x)a-4ax))$
$=x+4\varepsilon(-(v,x)a-(a,x)v-vx-4v(ax)))$
(2) $=x+8\varepsilon[L_{a},L_{v}](x).$
Now Proposition 5.2 shows $[L_{a},L_{v}]$ is a derivation. For each of these cases, $\langle\langle a,b\rangle\rangle_{\frac{1}{2}}(a)=\langle v\rangle$. This implies, by the Seress Lemma (Lemma 2.4), that $[L_{a},L_{b}]=\mu[L_{a},L_{v}]$ for a certain $\mu\in\mathbb{F}$. This means $[L_{a},L_{b}]$ is also a derivation. ∎
We now prove a converse to Corollary 5.3. When the characteristic is zero, this is easily done by taking the exponential of the derivations. In positive characteristic however, this is harder. For that reason, we will prove this using a different technique.
Lemma 5.4.
Suppose $P\in A[\lambda]$ is a polynomial such that $P(\mu+\nu\varepsilon)=0\in A\otimes_{\mathbb{F}}\mathbb{F}[\varepsilon]$ for certain $\mu\in\mathbb{F},\nu\in\mathbb{F}^{\times}$. Then $\mu$ is a root of multiplicity at least $2$ of $P$.
Proof.
Write $P(\lambda)=a_{n}\lambda^{n}+a_{n-1}\lambda^{n-1}\dots+a_{0}$. Since $\varepsilon^{2}=0$, we can compute that $P(\mu+\nu\varepsilon)=P(\mu)+\nu\varepsilon(na_{n}\mu^{n-1}+(n-1)a_{n-1}\mu^{n-2}+\dots a_{1})=P(\mu)+\nu\varepsilon P^{\prime}(\mu)$, where $P^{\prime}$ denotes the formal derivative of $P$. Since $P(\mu+\nu\varepsilon)$ has to be zero, this implies $\mu$ is a root for both $P$ and its formal derivative. This implies $\mu$ is a root of multiplicity at least $2$. ∎
Proposition 5.5.
Let $(A,X)$ be a primitive axial algebra of Jordan type half over a field $\mathbb{F}$ with $\operatorname{char}\mathbb{F}\neq 3$. If $a,b\in X$ such that $D_{a,b}$ is a derivation, then $\langle\langle a,b\rangle\rangle$ is solid.
Proof.
Clearly, we only have to prove this when $(a,b)=\frac{1}{4}$ and $\langle\langle a,b\rangle\rangle$ is $3$-dimensional. We may assume without loss of generality that $\mathbb{F}$ is algebraically closed. Note that $\rho=\tau_{a}\tau_{b}$ has order $3$, since $(\tau_{a}\tau_{b})^{3}=\tau_{a^{\tau_{b}}}\tau_{b^{\tau_{a}}}=0$.
Note that $\chi=\tau_{a}\tau_{a+\varepsilon v}=\operatorname{id}_{A_{\mathbb{F}[\varepsilon]}}+8\varepsilon[L_{a},L_{v}]$ is an automorphism of $A_{\mathbb{F}[\varepsilon]}$ by Propositions 5.2 and 2, where $v=e-f$. This means that if $a_{\lambda}$ is an axis of $A_{\mathbb{F}[\varepsilon]}$, then $a_{\lambda}^{\chi}=a_{\lambda(1-2\varepsilon)}$ is also an axis of $A_{\mathbb{F}[\varepsilon]}$. So we have $\lambda^{2}Q_{x}^{t}(\lambda(1-2\varepsilon))=\lambda^{2}P_{x,y}^{t}(\lambda(1-2\varepsilon))=0$ for every $x,y\in A$, for at least $3$ different values of $\lambda\in\mathbb{F}^{\times}$. This means, by Lemma 5.4, that the polynomials $\lambda^{2}Q_{x}^{t},\lambda^{2}P_{x,y}^{t}\in A[\lambda]$ have at least 6 roots (counted with multiplicity) for every $x,y\in A$. Since $\deg_{\lambda}\lambda^{2}Q_{x}^{t},\deg_{\lambda}\lambda^{2}P_{x,y}^{t}\leq 4$, this means $\lambda^{2}Q_{x}^{t},\lambda^{2}P_{x,y}^{t}=0$.
The previous paragraph implies that $Q_{x}^{t}(\lambda)=P_{x,y}^{t}(\lambda)=0$ for every $x,y\in A$ and $\lambda\in\mathbb{F}^{\times}$, and Lemma 4.1 in turn shows $\langle\langle a,b\rangle\rangle$ is solid. ∎
Proposition 5.6.
Let $(A,X)$ be a primitive axial algebra of Jordan type half over a field $\mathbb{F}$ with $\operatorname{char}\mathbb{F}=3$. If $a,b\in X$ such that $D_{a,b}$ is a derivation, then $\langle\langle a,b\rangle\rangle$ is solid.
Proof.
Clearly, we only have to prove this when $(a,b)=1$ and $\langle\langle a,b\rangle\rangle$ is $3$-dimensional. Note that $\rho=\tau_{a}\tau_{b}$ has order $3$, since $(\tau_{a}\tau_{b})^{3}=\tau_{a^{\tau_{b}}}\tau_{b^{\tau_{a}}}=0$. Note that in the unipotent case, $P^{u}_{x,y}$ and $Q^{u}_{x}$ are always polynomial maps.
Note that $\chi=\tau_{a}\tau_{a+\varepsilon v}=\operatorname{id}_{A_{\mathbb{F}[\varepsilon]}}+8\varepsilon[L_{a},L_{v}]$ is an automorphism of $A_{\mathbb{F}[\varepsilon]}$ by Propositions 5.2 and 2, where $v=ab$ when $(a,b)=0$ and $v=2(ab-a)$ when $(a,b)=1$. This means that if $a_{\lambda}$ is an axis, then $a_{\lambda}^{\chi}=a_{\lambda-4\varepsilon}$ is also an axis. This means $Q_{x}^{u}(\lambda-4\varepsilon)=P_{x,y}^{u}(\lambda-4\varepsilon)=0$ for every $x,y\in A$ and every $\lambda\in\mathbb{F}_{3}$. This means, by Lemma 5.4, that the polynomials $Q_{x}^{u},P_{x,y}^{u}\in A[\lambda]$ have at least 6 roots (counted with multiplicity) for every $x,y\in A$. Since $\deg_{\lambda}Q_{x}^{u},\deg_{\lambda}P_{x,y}^{u}\leq 4$.
The previous paragraph implies $Q_{x}^{u}=P_{x,y}^{u}=0$ for every $x,y\in A$, and Lemma 4.1 shows $\langle\langle a,b\rangle\rangle$ is solid. ∎
Theorem 5.7.
Let $(A,X)$ be a primitive axial algebra of Jordan type half over a field $\mathbb{F}$ with $\operatorname{char}\mathbb{F}\neq 2$. For any $a,b\in X$, the line $\langle\langle a,b\rangle\rangle$ is solid if and only if $D_{a,b}$ is a derivation.
Proof.
This follows from Corollaries 5.3, 5.5 and 5.6. ∎
6. Proof of Theorem 1.2
First we need the definition of an almost Jordan ring.
Definition 6.1.
A non-associative ring $R$ is almost Jordan if for all $x,y\in R$ the identity
(3) $2(yx\cdot x)\cdot x+yx^{3}=3(yx^{2})x$
holds.
Such algebras and their relation to Jordan algebras were first studied by Osborn in [16], but the name was given much later, in [10].
Lemma 6.2.
A non-associative commutative ring $R$ with $2\in R^{\times}$ is almost Jordan if and only if $D_{x,y}$ is a derivation of $R$ for all $x,y\in R$.
Proof.
This is proven in [16, p. 1115, Equations (4) and (5)]. ∎
Corollary 6.3.
Suppose $(A,X)$ is a primitive axial algebra of Jordan type half, and $X$ spans $A$ linearly. If for every $a,b\in X$ the line $\langle\langle a,b\rangle\rangle$ is solid, then $A$ is almost Jordan.
Proof.
For any $a,b\in X$, the map $[L_{a},L_{b}]$ is a derivation by Theorem 5.7. Since $X$ linearly spans $A$, this implies $[L_{x},L_{y}]$ is a derivation for any $x,y\in A$. Lemma 6.2 then shows $A$ is almost Jordan. ∎
As was noted in [16], the existence of idempotents very often forces almost Jordan algebras to be Jordan. We prove a similar result to the results in [16] that is applicable to our situation.
Proposition 6.4.
An almost Jordan algebra $A$ over a field $\mathbb{F}$ with $\operatorname{char}\mathbb{F}\neq 2$ that is linearly spanned by idempotents is a Jordan algebra.
Proof.
First off, by [16, Equation (6)], if $a\in A$ is any idempotent, its multiplication operator $\operatorname{ad}(a)$ is diagonalizable with eigenvalues $0,1$ and $\frac{1}{2}$. If we linearize the identity $x^{2}(ax)-x(ax^{2})$ and divide the resulting identity by $2$, we get
(4) $(yz)(ax)+(xy)(az)+(xz)(ay)-((yz)a)x-((xy)a)z-((xz)a)y.$
Because $A$ is linearly spanned by axes, we may assume $a$ is an axis, and $x,y,z$ are eigenvectors of $\operatorname{ad}(a)$ with eigenvalue $\lambda_{x},\lambda_{y},\lambda_{z}\in\{1,0,\tfrac{1}{2}\}$. Then Equation 4 becomes
(5) $\lambda_{x}(yz)x+\lambda_{z}(xy)z+\lambda_{y}(xz)y-((yz)a)x-((xy)a)z-((xz)a)y.$
Because of symmetry, we need to check this for $(\lambda_{x},\lambda_{y},\lambda_{z})=(1,1,1)$, $(1,1,\frac{1}{2})$, $(1,1,0)$, $(1,\frac{1}{2},\frac{1}{2})$, $(1,\frac{1}{2},0)$, $(1,0,0)$, $(\frac{1}{2},\frac{1}{2},\frac{1}{2})$, $(\frac{1}{2},\frac{1}{2},0)$, $(\frac{1}{2},0,0)$ and $(0,0,0)$. The cases $(1,1,1)$, $(0,0,0)$ and $(1,0,0)$ are easily checked, and they hold in every primitive axial algebra. The same is true for $(1,1,\frac{1}{2})$ and $(1,1,0)$.
The case $(\lambda_{x},\lambda_{y},\lambda_{z})=(1,\frac{1}{2},\frac{1}{2})$.
In this case, we may suppose $x=a$. Then Equation 4 becomes
$(yz)a+\frac{1}{4}yz+\frac{1}{4}yz-((yz)a)a-\frac{1}{4}yz-\frac{1}{4}yz.$
This is zero, since $(yz)a$ is equal to $a$ up to a scalar, so $((yz)a)a=(yz)a$.
The case $(\lambda_{x},\lambda_{y},\lambda_{z})=(1,\frac{1}{2},0)$.
Again supposing $x=a$, Equation 4 becomes
$(yz)a-((yz)a)a-((ay)a)z$ $=\frac{1}{2}yz-\frac{1}{4}yz-\frac{1}{4}yz=0.$
The case $(\lambda_{x},\lambda_{y},\lambda_{z})=(\frac{1}{2},\frac{1}{2},0)$.
In this case, Equation 4 becomes
$\frac{1}{2}(yz)x+\frac{1}{2}(xz)y-\frac{1}{2}(yz)x-\frac{1}{2}(xz)y=0.$
The case $(\lambda_{x},\lambda_{y},\lambda_{z})=(\frac{1}{2},\frac{1}{2},\frac{1}{2})$.
Since $A$ is almost Jordan, we have
$[L_{a},L_{x}](yz)=[L_{a},L_{x}](y)z+[L_{a},L_{x}](z)y,$
which, when written out, becomes
$(a(yz))x-\frac{1}{2}(yz)x=\frac{1}{2}(xy)z+\frac{1}{2}(xz)y-(a(xy))z-(a(xz))y.$
But this is equivalent to Equation 4 being zero for $(\lambda_{x},\lambda_{y},\lambda_{z})=(\frac{1}{2},\frac{1}{2},\frac{1}{2})$.
The case $(\lambda_{x},\lambda_{y},\lambda_{z})=(\frac{1}{2},0,0)$.
In this case, Equation 4 becomes
$\frac{1}{2}(yz)x-\frac{1}{2}(xy)z-\frac{1}{2}(xz)y,$
which is zero by [16, Lemma 2, Equation (9)].
This implies $A$ satisfies the linearized Jordan identity. If $\operatorname{char}\mathbb{F}\neq 3$, then this is sufficient for $A$ to be Jordan. If $\operatorname{char}\mathbb{F}=3$, then [18, Lemma 3.1] shows $A$ is Jordan. ∎
This is enough to show that a primitive axial algebra of Jordan type half $(A,X)$ such that $\langle\langle a,b\rangle\rangle$ is solid for all $a,b\in X$ and $X$ spans $A$ linearly, is actually a Jordan algebra. By using the maps $D_{a,b}$ and Theorem 5.7, we can prove more. The ideas in the proof of the following lemma are not from the author, they are largely adapted from a conversation with Sergey Shpectorov.
Lemma 6.5.
Let $(A,X)$ be a primitive axial algebra of Jordan type half over $\mathbb{F}$ with $\operatorname{char}\mathbb{F}\neq 2$. If $a,b,c\in X$ are such that $\langle\langle a,b\rangle\rangle$ and $\langle\langle a,c\rangle\rangle$ are solid, then $\langle\langle a,c^{\tau_{b}}\rangle\rangle$ is also solid. Moreover, if not $(a,b)=(a,c)=(a,bc)=0$, then $\langle\langle b,c\rangle\rangle$ is also solid.
Proof.
We may assume $\mathbb{F}$ is algebraically closed. Suppose first that $\langle\langle a,b\rangle\rangle$ is not flat. Then we can choose $p_{1},p_{2}\in\langle\langle a,c\rangle\rangle$ such that $(a,p_{1}),(a,p_{2})\neq\frac{1}{4}$ and $a,p_{1},p_{2}$ span $\langle\langle a,c\rangle\rangle$ linearly. Since $\langle\langle a,b\rangle\rangle$ is not flat, the variety of axes $V$ in $\langle\langle a,b\rangle\rangle$ is irreducible. This implies that the maps $\pi_{i}\colon V\to\mathbb{F}\colon q\mapsto(p_{i},q)$ are either constant or dominant, and thus there are only finitely many axes $q\in V$ with $(q,p_{1})=\frac{1}{4}$ or $(q,p_{2})=\frac{1}{4}$. But then we can find $q_{1},q_{2}\in V$ such that $a,q_{1},q_{2}$ span $\langle\langle a,b\rangle\rangle$ linearly and $(p_{i},q_{j})\neq\frac{1}{4}$. This implies $[L_{p_{i}},L_{q_{j}}]$ is a derivation of the algebra $A$ for $i,j\in\{1,2\}$, and so are $[L_{a},L_{q_{i}}],[L_{a},L_{p_{i}}]$. This implies $[L_{x},L_{y}]$ is a derivation for every $x\in\langle\langle a,b\rangle\rangle,y\in\langle\langle a,c\rangle\rangle$, and thus also $[L_{b},L_{c}]$. Thus $\langle\langle b,c\rangle\rangle$ is solid.
If $(a,b)=(a,c)=0$ but $(a,bc)=(c,ab)\neq 0$, then by a similar reasing there exist $\lambda,\mu\in\mathbb{F}^{\times}$ such that $\langle\langle a+\lambda ab,c\rangle\rangle,\langle\langle b+\mu ab,c\rangle\rangle$ are solid, and thus $[L_{a},L_{c}],[L_{a+\lambda a\cdot b},L_{c}],[L_{b+\mu a\cdot b},L_{c}]$ are derivations. But then $[L_{b},L_{c}]$ is a derivation, thus $\langle\langle b,c\rangle\rangle$ is solid by Theorem 5.7.
In either one of these cases, $\langle\langle a,b\rangle\rangle$, $\langle\langle a,c\rangle\rangle$ and $\langle\langle b,c\rangle\rangle$ are solid. We know in either one of these cases that for any $x\in\langle\langle a,b\rangle\rangle$ both $D_{x,a}$ and $D_{x,c}$ are derivations. This implies that $D_{a^{\tau_{b}},c}$ is a derivation, so $D_{a,c^{\tau_{b}}}$ is as well.
Now suppose $(a,b)=(a,c)=(a,bc)=0$ and $(b,c)=\frac{1}{4}$. Then $(a,c^{b})=(a,b+c-bc)=0$, and thus $\langle\langle a,b^{c}\rangle\rangle$ is solid by Theorem 5.7. ∎
Theorem 6.6.
A primitive axial algebra of Jordan type half $(A,X)$ of Jordan type half over a field $\mathbb{F}$ with $\operatorname{char}\mathbb{F}\neq 2$ such that for all $a,b\in X$ the line $\langle\langle a,b\rangle\rangle$ is solid, is Jordan.
Proof.
By [6, Corollary 1.2], the set $\overline{X}=X^{\operatorname{Miy}(X)}$ spans the axial algebra $A$ linearly. If we prove that for all $a,b\in\overline{X}$ that the subalgebra $\langle\langle a,b\rangle\rangle$ is solid, the theorem follows from Corollaries 6.3 and 6.4. We prove that for $a,b\in X$ and $g,h\in\operatorname{Miy}(X)$ that $\langle\langle a^{g},b^{h}\rangle\rangle$ is solid by induction on the length of $gh^{-1}$ in terms of the generators $\tau_{a},a\in X$. Indeed, let $a,b,c\in X$. We need to prove the line $a,b^{\tau_{c}}$ is solid. But this follows immediately from Lemma 6.5. Now, the line $\langle\langle a^{\tau_{c_{1}}\dots\tau_{c_{k}}},b^{\tau_{c_{n}}\dots\tau_{c_{k+1}}}\rangle\rangle$ is solid if and only if $\langle\langle a^{\tau_{c_{1}}\dots\tau_{c_{n}}},b\rangle\rangle$ is. By induction, $\langle\langle a^{\tau_{c_{1}}\dots\tau_{c_{n-1}}},b\rangle\rangle$ and $\langle\langle a^{\tau_{c_{1}}\dots\tau_{c_{n-1}}},c_{n}\rangle\rangle$ are solid, so by Lemma 6.5, so is $\langle\langle a^{\tau_{c_{1}}\dots\tau_{c_{n-1}}},b^{\tau_{c_{n}}}\rangle\rangle$. But this is equivalent to $\langle\langle a^{\tau_{c_{1}}\dots\tau_{c_{n}}},b\rangle\rangle$ being solid. ∎